A rich aunt has promised you $2,090.00, one year from today. In addition, each year after that , she has promised a payment (on the anniversary of the last payment) that is 2.00% larger than the last payment. She will continue to show this generosity for 25 years, giving a total of 25 payments. If the interest rate is 8.00%, what is her promise worth today?
Answer #1

This can be soldu using formula of PV of growing annuity
PV of growing annuity = P/r-g[1-((1+g)/(1+r))^n]
Here P = first payment = 2090$
r = rate of interest = 8%
g= Growth rate = 2%
n = no of payments = 25
Pv of growing annuity = 2090/(8%-2%) [ 1 -((1+2%)/(1+8%))^25]
=2090/6%[1-((1.02)/1.08))^25]
=34833.33[1-0.239558]
=26488.73$
Thus her promise is worth $ 26488.73 today